# Space Group Informed Transformer for Crystalline Materials Generation

### v2

@@ Abstract
We introduce CrystalFormer, a transformer-based autoregressive model specifically designed for space group-controlled generation of crystalline materials. The incorporation of space group symmetry significantly simplifies the crystal space, which is crucial for data and compute efficient generative modeling of crystalline materials. Leveraging the prominent discrete and sequential nature of the Wyckoff positions, CrystalFormer learns to generate crystals by directly predicting the species and locations of symmetry-inequivalent atoms in the unit cell. We demonstrate the advantages of CrystalFormer in standard tasks such as symmetric structure initialization and element substitution compared to conventional methods implemented in popular crystal structure prediction software. Moreover, we showcase the application of CrystalFormer of property-guided materials design in a plug-and-play manner. Our analysis shows that CrystalFormer ingests sensible solid-state chemistry knowledge and heuristics by compressing the material dataset, thus enabling systematic exploration of crystalline materials. The simplicity, generality, and flexibility of CrystalFormer position it as a promising architecture to be the foundational model of the entire crystalline materials space, heralding a new era in materials modeling and discovery.
@@ I Introduction
It is anticipated that generative model-based approaches will introduce groundbreaking changes to the traditional workflows of material discovery. A generative pre-trained foundation model for crystalline materials is a key step towards such a lofty goal. However, despite intensive efforts [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], the current generative models for crystalline materials fall short to match the success of other domains. Simply scaling the compute and model size of the current crystal generative model may not be feasible because the amount of high-quality data for crystalline materials is much less than compared to language and image domains. Therefore, leveraging the inherent inductive biases specific to crystalline structures for more data-efficient generative modeling is essential, as has been pursued in some of recent works [23, 24, 25, 26].
@@
Table 1: A back-of-envelope estimate of the size of the crystalline material space. In the ”$P1$ world”, one treats crystals as if they were in the first and the least symmetric $P1$ space group. For the estimate, we consider 100 possible chemical elements and 20 atoms in the unit cell with a coordinate grid size of 100 in each direction. In the case of utilizing the symmetry of a typical space group, we consider 5 symmetry inequivalent atoms occupying 10 possible Wyckoff positions. The additional factor of 100 accounts for the remaining degree of freedom for the fractional coordinates. See Refs. [27, 28] for alternate estimates of the materials space in the context of crystal structure prediction.
@@
Figure 1: (a) The Wyckoff positions of the $R\bar{3}c$ space group (No. 167). We highlight the occupied Wyckoff positions of calcite CaCO3 crystal which belongs to this space group. Carbon, calcium, and oxygen atoms occupy the ’6a’, ’6b’, and ’18e’ positions, respectively. (b) The CrystalFormer is a decoder-only autoregressive transformer that models the space group controlled crystal structures by predicting probabilities of the Wyckoff letter $w_{i}$, chemical element $a_{i}$, and fractional coordinates $(x_{i},y_{i},z_{i})$ of each symmetry inequivalent atom, and finally, the lattice parametrized by $\boldsymbol{\ell}$ sequentially. (c) The crystal data of CaCO3 is summarized in a table. In the table, the x-coordinate of oxygen atom $x_{3}=0.257$ is the only continuous variable that needs to be predicted. All other fractional coordinates are fixed by discrete data like the space group number and Wyckoff letters. The string below the table shows the sequential representation of the CaCO3 crystal with space group, Wyckoff letter, and atom species as the input to the CrystalFormer model. (d) Autoregressive generation of the crystal. One first places carbon atoms at the ’6a’ position, then places calcium atoms at ’6b’ position, and finally places oxygen atoms at ’18e’ position. In each step of the sampling procedure, there is a choice of the Wyckoff positions, atom species, and the fractional coordinates if they are still unspecified.
@@
Space group symmetry imposes significant constraints on a crystal. First of all, the space group identifies the crystal system to which a crystal belongs, thereby limiting the permissible values for the lattice parameters that define the length and angles of the crystal’s unit cell. Moreover, the symmetry operations associated with a given space group ensure that symmetry equivalent atoms are consistently mapped among themselves in the crystal. This requirement enforces strict conditions regarding the types of chemical elements present, their specific locations within the crystal, and the number of each chemical species in the unit cell. A key concept to express these constraints is the Wyckoff positions, which delineate unique areas within a unit cell that are defined by the symmetry operations of the crystal’s space group. These positions are represented as fractional coordinates, enabling precise definition relative to the unit cell’s axes. For example, Fig. 1(a) shows the Wyckoff positions for the space group $R\bar{3}c$ (No. 167). The Wyckoff positions are labeled by letters in the alphabet, starting from special points in the bottom to general positions in the top. The multiplicity counts the number of equivalent positions connected by the space group symmetry operations. All of them should be occupied by the same type of atoms to uphold the space group symmetry. For example, the top row of the table in Fig. 1(a) contains the general position $(x,y,z)$ that can be mapped to 36 positions under the symmetry operations of the $R\bar{3}c$ space group.
@@
Nature tends to place atoms in those special Wyckoff positions at the bottom of the table. For example, we highlight the occupied Wyckoff positions of calcite (CaCO3) crystal in Fig. 1, associated with the $R\bar{3}c$ space group. One sees that the Wyckoff letter ’6a’ and ’6b’ deterministically define the locations of the carbon and calcium atoms within the unit cell. In addition, it follows that $a=b$, and $\alpha=\beta=90^{\circ},\gamma=120^{\circ}$ as the $R\bar{3}c$ space group belongs to the trigonal crystal system. Ultimately, despite having 30 atoms in the unit cell, there are only three continuous degrees of freedom for the CaCO3 structure: the x-coordinate of oxygen atom $0.257$ and the lattice constants $a=b=4.99\AA$ and $c=17.07\AA$. All other information about the crystal structure can be specified via discrete data such as the Wyckoff letters and chemical species.
@@
The prominent discrete and sequential features illustrated in Figure 1 are ubiquitous in crystalline materials. The Wyckoff positions not only specify possible locations of atoms in the unit cell, but their associated multiplicities also put strong constraints on the number of atoms. Therefore, space group symmetry significantly reduces the degrees of freedom of crystalline materials. Failing to exploit this information in generative modeling not only renders learning inefficient, it also severely impairs the generalization ability of the model. For example, the performance of the generative model quickly deteriorates as the number of atoms increases due to it is challenging to generate highly symmetric crystal structures [16]. On the other hand, statistical analysis shows that the Wyckoff sequences of known inorganic compounds  [33] are far from being exhausted, implying there are statistical correlations to be exploited to compress the materials database.
@@
The organization of the present paper is as follows. Section II introduces the model architecture, training, and sampling of the CrystalFormer. Section III reveals the chemical intuition encoded in the trained model by inspecting its weights and generated crystal samples. Such inspection also builds up an understanding of the strength of the model. Section IV demonstrates applications CrystalFormer for symmetric structure initialization and element substitution, where it exhibits great efficiency and generality compared to existing approaches. Furthermore, we showcase CrystalFormer’s ability for property-guided materials design in a plug-and-play manner. Section V puts the key contribution of present work in the broad context of crystal generative models. Finally, Section VI outlooks for possible extension and future of the present line of research. We have released the codes and trained model at [34].
II CrystalFormer
@@
II.1 Model
To exploit the space group symmetry of the crystal, we focus on the Wyckoff positions of symmetry-inequivalent atoms. Wyckoff letters follow the alphabetical order, where ”a” stands for the positions with the highest order of symmetry for the given space group. Later letters in the alphabet indicate more general positions with reduced site symmetries. Note that the information of the space group number and Wyckoff letter fully determine the multiplicities. In cases where the atom positions are not fully fixed by the Wyckoff letter, we will also consider the remaining fractional coordinates, e.g. the $x$-coordinate of the oxygen atoms in the CaCO3 example shown in Fig. 1. To generate crystals, one samples the Wyckoff letter, chemical element, and fractional coordinates of each atom sequentially. The sampling procedure starts from special higher symmetry sites with smaller multiplicities and then goes on to general lower symmetry regions with larger multiplicities.
With these considerations, we define a crystal data as $\boldsymbol{\mathcal{C}}=\{\boldsymbol{W},\boldsymbol{A},\boldsymbol{X},\boldsymbol{L}\}$. Here $\boldsymbol{W}=[w_{1},w_{2},\ldots,w_{n}]$ are Wyckoff letters and $\boldsymbol{A}=[a_{1},a_{2},\ldots,a_{n}]$ are chemical species. Here, $n$ stands for the number of symmetrically inequivalent atoms in the conventional unit cell. For example, as shown in Fig. 1(b) one has $n=3$ for CaCO3. Explicitly including the Wyckoff letter in the generative modeling is the key of the present work. Next, $\boldsymbol{X}=[(x_{i},y_{i},z_{i})]\in\mathbb{R}^{n\times 3}$ are the fractional coordinates of symmetrically inequivalent atoms. Lastly, $\boldsymbol{L}=[a,b,c,\alpha,\beta,\gamma]$ denotes the lattice parameters of the conventional unit cell of the crystal.
@@ II.1 Model
At first sight, it may appear unnatural to employ an autoregressive model for crystals since there seems to be no obvious order for atoms in the unit cell. However, the sequential nature of Wyckoff positions suggests a natural way to arrange symmetrically inequivalent atoms in an alphabetical order of the Wyckoff letters. Following this key observation, we represent crystal data as sequences of space groups, Wyckoff letters, chemical species, and fractional coordinates of each symmetrically inequivalent atom. Together with the information lattice parameters, such sequence fully characterizes the compositional and structural information of crystalline material. Since statistical analysis reveals that anions are in less symmetric positions than cations for inorganic crystals [30], one would expect that anion atoms will typically appear after cation atoms in such a sequence. For example, CaCO3 is represented as a string ”167-a-C-0-0-1/4-b-Ca-0-0-0-e-O-0.257-0-1/4”. Autoregressive sampling of such a string means the model generates the crystal by placing the atoms sequentially into the unit cell, starting from the special position with high site symmetry to the general position with the lowest site symmetry, see Fig. 1(d).
@@
We build CrystalFormer, an autoregressive transformer [40] to model the space group conditioned-probability distribution of crystalline materials Eq. (1). The space group number $g$ is the first input to CrystalFormer. The remaining inputs are the Wyckoff letter, chemical species, and fractional coordinates of each atom. One can go through the table of Fig. 1(b) in a raster order to collect these atomistic features. We feed vector embeddings of the space group number, Wyckoff letter, and the chemical species input to the CrystalFormer. In particular, we also concatenate the vector embedding of $g$ to all other inputs since it is the key control variable for the crystal generation. Moreover, we have also provided the multiplicity of each Wyckoff position as an additional feature. The multiplicity can be easily inferred from the space group and the Wyckoff letters. We feed the fractional coordinates as Fourier features into the transformer so that the model preserves the periodicity of the unit cell [41, 13]. We pad the atom sequence up to a maximum length and treat the output as parameters of the conditional probability distribution Eq. (1), see Fig. 1(b). At the location of the first padding atom, we predict the lattice parameters.
@@
We implement a number of constraints in the model to further reduce its phase space. First, the Wyckoff letters should be valid for the given space group. For example, for the space group $R\bar{3}c$ (No. 167) the Wyckoff letters go from ’a’ up to ’f’. Second, we require that the Wyckoff letters $w_{i}$ follow alphabetical order in the sequence  [^1] We follow the convention that capital letters appear after lower case letters. This handles the edge case of the $Pmmm$ space group (No. 47) whose Wyckoff letters used up 26 lowercase letters and reached ’A’ for the generic position. In addition, we use the letter ’X’ to indicate the Wyckoff position of padding atoms that appear at the end of the sequence, see e.g. $w_{4}$ of Fig. 1(b).. Lastly, the Wyckoff positions with no free fractional coordinates (such as ’a’, ’b’, and ’d’ positions in the $R\bar{3}c$ space group) can only be occupied once. Those constraints are implemented by setting the logit biases of Wyckoff letters to mask out invalid sequences  [43, 44].
@@
The design of CrystalFormer focuses mostly on the space group symmetries which we believe to be the most important inductive bias for crystalline materials. This design decision significantly impacts the treatment of other symmetries. First, it is often possible to place the origin of the unit cell at the inversion center of the specified space group. The chosen origin naturally fixes the continuous translation invariance of fractional coordinates. Second, by only considering symmetry-inequivalent atoms and labeling them with Wyckoff letters, one fixes most of the permutation invariance in the representation. For those Wyckoff positions with continuous degrees of freedom, there may be multiple symmetry-inequivalent atoms with the same Wyckoff letters. We arrange these atoms according to the lexicographic order of fractional coordinates [45] in the sequence. Note that in a crystal environment, the same type of atoms occupying different Wyckoff positions could be regarded as distinguished particles as they generally have different site symmetry. Lastly, the periodicity of the fractional coordinates is respected in CrystalFormer since they are treated as periodic variables under the von Mises distribution.
II.2 Training
The CrystalFormer is trained by minimizing the negative log-likelihood over training dataset
$\mathcal{L}=-\mathop{\mathbb{E}}_{\boldsymbol{\mathcal{C}},g}\left[\ln p(\boldsymbol{\mathcal{C}}|g)\right].$ (2)
@@ II.2 Training
Writing out $p(\boldsymbol{\mathcal{C}}|g)$ according to Eq. (1), the objective function contains the negative log-likelihood of discrete variables such as Wyckoff letters $\boldsymbol{W}$ and chemical species $\boldsymbol{A}$, as well as continuous variables such as fractional coordinates $\boldsymbol{X}$ and the lattice parameters $\boldsymbol{L}$. In the objective function, for continuous variables $\boldsymbol{X},\boldsymbol{L}$ we consider only active ones that are not fixed by the space groups and Wyckoff letters. In this way, those special fractional coordinates (e.g. 00, $\frac{1}{4}$) and lattice parameters (e.g. $90^{\circ}$, $120^{\circ}$ ) do not contribute to the loss function.
@@
In the present work, we train the CrystalFormer using the MP-20 dataset [11] which is a popular dataset that represents a majority of experimentally known crystalline materials at ambient conditions with no more than 20 atoms in the primitive unit cell. The training dataset contains 27136 crystal structures. The subdivision of the training samples according to the space group has greatly reduced the number of samples in each space group category. On top of that, the distribution of training samples is quite uneven among the space groups, which reflects the imbalanced distribution of crystals over space groups in nature [30]. In fact, there is no training data in 61 out of 230 space groups as shown in Fig. S2. Nevertheless, we still employ the MP-20 as the training set so that the performance of the model can be more easily gauged with the others in the literature, see appendix B. Figure 2 shows a breakup of the learning curves for the Wyckoff position, chemical species, fractional coordinates, and lattice parameters. We select the model checkpoint with the lowest total validation loss to generate crystal samples.

### v3

@@ Abstract
We introduce CrystalFormer, a transformer-based autoregressive model specifically designed for space group-controlled generation of crystalline materials. By explicitly incorporating space group symmetry, CrystalFormer greatly reduces the effective complexity of crystal space, which is essential for data-and compute-efficient generative modeling of crystalline materials. Leveraging the prominent discrete and sequential nature of the Wyckoff positions, CrystalFormer learns to generate crystals by directly predicting the species and coordinates of symmetry-inequivalent atoms in the unit cell. We demonstrate the advantages of CrystalFormer in standard tasks such as symmetric structure initialization and element substitution over widely used conventional approaches. Furthermore, we showcase its plug-and-play application to property-guided materials design, highlighting its flexibility. Our analysis reveals that CrystalFormer ingests sensible solid-state chemistry knowledge and heuristics by compressing the material dataset, thus enabling systematic exploration of crystalline materials space. The simplicity, generality, and adaptability of CrystalFormer position it as a promising architecture to be the foundational model of the entire crystalline materials space, heralding a new era in materials discovery and design.
Keywords: inorganic crystals, generative model, autoregressive transformer, space group symmetry
(Received 10 February 2025; Revised 5 May 2025; Accepted 12 September 2025)
@@ I Introduction
It is anticipated that generative model-based approaches will introduce groundbreaking changes to the traditional workflows of material discovery. A generative pre-trained foundation model for crystalline materials is a key step towards such a lofty goal. However, despite intensive efforts [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], the current generative models for crystalline materials fall short to match the success of other domains. Simply scaling the compute and model size of the current crystal generative model may not be feasible because the amount of high-quality data for crystalline materials is much less compared to language and image domains. Therefore, leveraging the inherent inductive biases specific to crystalline structures for more data-efficient generative modeling is essential, as has been pursued in some of recent works [23, 24, 25, 26].
@@
Table 1: A back-of-envelope estimate of the size of the crystalline material space. In the “$P1$ world”, one treats crystals as if they were in the first and the least symmetric $P1$ space group. For the estimate, we consider 100 possible chemical elements and 20 atoms in the unit cell with a coordinate grid size of 100 in each direction. In the case of utilizing the symmetry of a typical space group, we consider 5 symmetry inequivalent atoms occupying 10 possible Wyckoff positions. The additional factor of 100 accounts for the remaining degree of freedom for the fractional coordinates and lattice parameters. See Refs. [27, 28] for alternate estimates of the materials space in the context of crystal structure prediction.
@@
[…]
@@
Space group symmetry imposes significant constraints on a crystal. First of all, the space group identifies the crystal system to which a crystal belongs, thereby limiting the permissible values for the lattice parameters that define the length and angles of the crystal’s unit cell. Moreover, the symmetry operations associated with a given space group ensure that symmetry equivalent atoms are consistently mapped among themselves in the crystal. This requirement enforces strict conditions regarding the types of chemical elements present, their specific locations within the crystal, and the number of each chemical species in the unit cell. A key concept to express these constraints is the Wyckoff positions, which delineate unique areas within a unit cell that are defined by the symmetry operations of the crystal’s space group. These positions are represented as fractional coordinates, enabling precise definition relative to the unit cell’s axes. For example, Fig. 1a shows the Wyckoff positions for the space group $R\bar{3}c$ (No. 167). The Wyckoff positions are labeled by letters in the alphabet, starting from special points in the bottom to general positions in the top. The multiplicity counts the number of equivalent positions connected by the space group symmetry operations. All of them should be occupied by the same type of atoms to uphold the space group symmetry. For example, the top row of the table in Fig. 1a contains the general position $(x,y,z)$ that can be mapped to 36 positions under the symmetry operations of the $R\bar{3}c$ space group.
@@
Nature tends to place atoms in those special Wyckoff positions at the bottom of the table. For example, we highlight the occupied Wyckoff positions of calcite (CaCO3) crystal in Fig. 1, associated with the $R\bar{3}c$ space group. One sees that the Wyckoff letter ‘6a’ and ‘6b’ deterministically define the locations of the carbon and calcium atoms within the unit cell. In addition, it follows that $a=b$, and $\alpha=\beta=90^{\circ},\gamma=120^{\circ}$ as the $R\bar{3}c$ space group belongs to the trigonal crystal system. Ultimately, despite having 30 atoms in the unit cell, there are only three continuous degrees of freedom for the CaCO3 structure: the x-coordinate of oxygen atom $0.257$ and the lattice constants $a=b=4.99\AA$ and $c=17.07\AA$ . All other information about the crystal structure can be specified via discrete data such as the Wyckoff letters and chemical species.
@@
The prominent discrete and sequential features illustrated in Fig. 1 are ubiquitous in crystalline materials. The Wyckoff positions not only specify possible locations of atoms in the unit cell, but their associated multiplicities also put strong constraints on the number of atoms. Therefore, space group symmetry significantly reduces the degrees of freedom of crystalline materials. Failing to exploit this information in generative modeling not only renders learning inefficient, it also severely impairs the generalization ability of the model. For example, the performance of the generative model quickly deteriorates as the number of atoms increases due to it is challenging to generate highly symmetric crystal structures [16]. On the other hand, statistical analysis shows that the Wyckoff sequences of known inorganic compounds  [33] are far from being exhausted, implying there are statistical correlations to be exploited to compress the materials database.
@@
II Method
We will first introduce the CrystalFormer model, then reveal the chemical intuition encoded in the trained model by inspecting generated crystal samples. These inspections also build up understandings of the strength of the model.
II.1 CrystalFormer
@@
II.1.1 Model
To exploit the space group symmetry of the crystal, we focus on the Wyckoff positions of symmetry-inequivalent atoms. Wyckoff letters follow the alphabetical order, where “a” stands for the positions with the highest order of site symmetry for the given space group. Later letters in the alphabet indicate more general positions with reduced site symmetries. Note that the information of the space group number and Wyckoff letter fully determine the multiplicities of the Wyckoff positions. In cases where the atom positions are not fully fixed by the Wyckoff letter, we will also consider the remaining fractional coordinates, e.g. the $x$-coordinate of the oxygen atoms in the CaCO3 example shown in Fig. 1. To generate crystals, one samples the Wyckoff letter, chemical element, and fractional coordinates of each atom sequentially. The sampling procedure starts from special higher symmetry sites with smaller multiplicities and then goes on to general lower symmetry sites with larger multiplicities.
With these considerations, we define a crystal data as $\boldsymbol{\mathcal{C}}=\{\boldsymbol{W},\boldsymbol{A},\boldsymbol{X},\boldsymbol{L}\}$. Here $\boldsymbol{W}=[w_{1},w_{2},\ldots,w_{n}]$ are Wyckoff letters and $\boldsymbol{A}=[a_{1},a_{2},\ldots,a_{n}]$ are chemical species. Here, $n$ stands for the number of symmetrically inequivalent atoms in the conventional unit cell. For example, as shown in Fig. 1b one has $n=3$ for CaCO3. Explicitly including the Wyckoff letter in the generative modeling is the key of the present work. Next, $\boldsymbol{X}=[(x_{i},y_{i},z_{i})]\in\mathbb{R}^{n\times 3}$ are the fractional coordinates of symmetrically inequivalent atoms. Lastly, $\boldsymbol{L}=[a,b,c,\alpha,\beta,\gamma]$ denotes the lattice parameters of the conventional unit cell of the crystal.
@@ II.1.1 Model
At first sight, it may appear unnatural to employ an autoregressive model for crystals since there seems to be no obvious order for atoms in the unit cell. However, the sequential nature of Wyckoff positions suggests a natural way to arrange symmetrically inequivalent atoms in an alphabetical order of the Wyckoff letters. Following this key observation, we represent crystal data as sequences of space groups, Wyckoff letters, chemical species, and fractional coordinates of each symmetrically inequivalent atom. Together with the information lattice parameters, such sequence fully characterizes the compositional and structural information of crystalline material. Since statistical analysis reveals that anions are in less symmetric positions than cations for inorganic crystals [30], one would expect that anion atoms will typically appear after cation atoms in such a sequence. For example, CaCO3 is represented as a string “167-a-C-0-0-1/4-b-Ca-0-0-0-e-O-0.257-0-1/4”. Autoregressive sampling of such a string means the model generates the crystal by placing the atoms sequentially into the unit cell, starting from the special position with high site symmetry to the general position with the lowest site symmetry, see Fig. 1d.
@@
We build CrystalFormer, an autoregressive transformer [39] to model the space group conditioned-probability distribution of crystalline materials Eq. (1). The space group number $g$ is the first input to CrystalFormer. The remaining inputs are the Wyckoff letter, chemical species, and fractional coordinates of each atom. One can go through the table of Fig. 1b in a raster order to collect these atomistic features. We feed vector embeddings of the space group number, Wyckoff letter, and the chemical species input to the CrystalFormer. In particular, we also concatenate the vector embedding of $g$ to all other inputs since it is the key control variable for the crystal generation. Moreover, we have also provided the multiplicity of each Wyckoff position as an additional feature. The multiplicity can be easily inferred from the space group and the Wyckoff letters. We feed the fractional coordinates as Fourier features into the transformer so that the model preserves the periodicity of the unit cell [40, 13]. We pad the atom sequence up to a maximum length and treat the output as parameters of the conditional probability distribution Eq. (1), see Fig. 1b. At the location of the first padding atom, we predict the lattice parameters.
@@
We implement a number of constraints in the model to further reduce its phase space. First, the Wyckoff letters should be valid for the given space group. For example, for the space group $R\bar{3}c$ (No. 167) the Wyckoff letters go from ‘a’ up to ‘f’. Second, we require that the Wyckoff letters $w_{i}$ follow alphabetical order in the sequence  [^1] We follow the convention that capital letters appear after lower case letters. This handles the edge case of the $Pmmm$ space group (No. 47) whose Wyckoff letters used up 26 lowercase letters and reached ‘A’ for the generic position. In addition, we use the letter ‘X’ to indicate the Wyckoff position of padding atoms that appear at the end of the sequence, see e.g. $w_{4}$ of Fig. 1b.. Lastly, the Wyckoff positions with no free fractional coordinates (such as ‘a’, ‘b’, and ‘d’ positions in the $R\bar{3}c$ space group) can only be occupied once. Those constraints are implemented by setting the logit biases of Wyckoff letters to mask out invalid sequences  [42, 43].
@@
The design of CrystalFormer focuses mostly on the space group symmetries which we believe to be the most important inductive bias for crystalline materials. This design decision significantly impacts the treatment of other symmetries. First, it is often possible to place the origin of the unit cell at the inversion center of the specified space group. The chosen origin naturally fixes the continuous translation invariance of fractional coordinates. Second, by only considering symmetry-inequivalent atoms and labeling them with Wyckoff letters, one fixes most of the permutation invariance over atom of the same type in the representation. For those Wyckoff positions with continuous degrees of freedom, there may be multiple symmetry-inequivalent atoms with the same Wyckoff letters. We arrange these atoms according to the lexicographic order of fractional coordinates [44] in the sequence. Note that in a crystal environment, the same type of atoms occupying different Wyckoff positions could be regarded as distinguished particles as they generally have different site symmetry. Lastly, the periodicity of the fractional coordinates is respected in CrystalFormer since they are treated as periodic variables following the von Mises distribution.
II.1.2 Training
The CrystalFormer is trained by minimizing the negative log-likelihood
$\mathcal{L}=-\mathop{\mathbb{E}}_{\boldsymbol{\mathcal{C}},g}\left[\ln p(\boldsymbol{\mathcal{C}}|g)\right],$ (2)
@@ II.1.2 Training
where the structures $\mathcal{C}$ and the corrosponding space group $g$ of crystals are sampled from the training dataset. Writing out $p(\boldsymbol{\mathcal{C}}|g)$ according to Eq. (1), the objective function contains the negative log-likelihood of discrete variables such as Wyckoff letters $\boldsymbol{W}$ and chemical species $\boldsymbol{A}$, as well as continuous variables such as fractional coordinates $\boldsymbol{X}$ and the lattice parameters $\boldsymbol{L}$. For the continuous variables $\boldsymbol{X},\boldsymbol{L}$ in the objective function, we consider only active ones that are not fixed by the space groups and Wyckoff letters. In this way, those special fractional coordinates (e.g. 0, $\frac{1}{4}$) and lattice parameters (e.g. $90^{\circ}$, $120^{\circ}$) which were already fixed by the chosen space group and Wyckoff letter will not not contribute to the loss function.
@@
In the present work, we train the CrystalFormer using the MP-20 dataset [11]. MP-20 is a popular dataset that represents a majority of experimentally known crystalline materials at ambient conditions with no more than 20 atoms in the primitive unit cell. The training dataset contains 27136 crystal structures. The subdivision of the training samples according to the space group has greatly reduced the number of samples in each space group category. On top of that, the distribution of training samples is quite uneven among the space groups, which reflects the imbalanced distribution of crystals over space groups in nature [30]. In fact, there is no training data in 61 out of 230 space groups as shown in Fig. S2. Nevertheless, we still employ the MP-20 as the training set so that the performance of the model can be more easily gauged with the others in the literature, see appendix B. Note that the CrystalFormer can generate reasonable samples even for those space groups without any training data. This because the model can exploit knowledge learned from other space groups to place suitable atoms in the Wyckoff positions due to weight sharing. Moreover, since the sampling process makes use the of Wyckoff position table. The three dimensional coordinates of atoms are not completely random even for unseen space groups. Fig. 2 shows a breakup of the learning curves for the Wyckoff position, chemical species, fractional coordinates, and lattice parameters. We select the model checkpoint with the lowest total validation loss to generate crystal samples.